OpenStudy (jiteshmeghwal9):

If \[\LARGE{\lim_{x \rightarrow 1}\frac{asin(x-1)+bcos(x-1)+4}{x^2-1}=-2}\], then a-b=??

Parth (parthkohli):

This is a zero-by-zero limit, so you can apply L'Hopital's Rule if you want to.

OpenStudy (jiteshmeghwal9):

\[\lim_{x \rightarrow 1}\frac{acos(x-1)-bsin(x-1)+4}{2x}\]am i correct??

Parth (parthkohli):

Wait, why did I call it a zero-by-zero limit -_-

OpenStudy (jiteshmeghwal9):

since the indeterminate form 0/0 is coming on putting x=1

Parth (parthkohli):

But for the limit to exist, yeah, the numerator must be zero. So \(b = -4\).

Parth (parthkohli):

And then you apply L'Hopital's rule.

OpenStudy (jiteshmeghwal9):

sorry but i don't gt why the numerator must be zero

OpenStudy (samigupta8):

In order for a limit to exist there shud be a zero in numerator otherwise the limit wouldnot be defined since there is a 0 in denominator

imqwerty (imqwerty):

yes since RHS is a finite quantity, at x->1 numerator must be 0 so 0+b+4=0 b=-4 and then try putting x=1+p and simplify

Parth (parthkohli):

^^^ To sum up, we need two conditions: 1. The limit should be 0/0-type. 2. The limit should be -2 using l'Hopital's rule.

OpenStudy (samigupta8):

a will also come out to be -4 so ur ans wud be 0

OpenStudy (jiteshmeghwal9):

thx :)